### 20 Jan 2024 v1

# The failure of Hölder regularity of solutions for the Camassa–Holm type equation in Besov spaces

Abstract: It is proved that if $u_{0}\in B^{s}_{p,r}$ with $s>1+\frac{1}{p},(p,r)\in[1,+\infty]\times[1,+\infty)$ or $s=1+\frac{1}{p},\ (p,r)\in[1,+\infty)\times\{1\}$, the solution of the Camassa–Holm equation belongs to $\mathcal{C}([0,T];B^{s}_{p,r})$. In the paper, we show that the continuity of the solution can not be improved to the Hölder continuity. Precisely speaking, the solution of the Camassa–Holm equation belongs to $\mathcal{C}([0,T];B^{s}_{p,r})$ but not to $\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with any $\alpha\in(0,1)$.
Keywords: Camassa–Holm equation, Hölder regularity, Besov spaces
MSC (2010): 35Q35; 35B30.
1 Introduction
In this paper, we are concerned with the Cauchy problem for the classical Camassa–Holm (CH) equation
$\begin{cases}u_{t}-u_{xxt}+3uu_{x}=2u_{x}u_{xx}+uu_{xxx},\;&(x,t)\in\mathbb{R}\times\mathbb{R}^{+},\\ u(x,t=0)=u_{0}(x),\;&x\in\mathbb{R}.\end{cases}$ (1.1)
Here the scalar function $u=u(t,x)$ stands for the fluid velocity at time $t\geq 0$ in the $x$ direction.
Setting $\Lambda^{-2}=(1-\partial^{2}_{xx})^{-1}$, then $\Lambda^{-2}f=G*f$ where $G(x)=\frac{1}{2}e^{-|x|}$ is the kernel of the operator $\Lambda^{-2}$. Thus, we can transform the CH equation (1.1) equivalently into the following transport type equation
$\begin{cases}\partial_{t}u+u\partial_{x}u=\mathbf{P}(u),\;&(x,t)\in\mathbb{R}\times\mathbb{R}^{+},\\ u(x,t=0)=u_{0}(x),\;&x\in\mathbb{R},\end{cases}$ (1.2)
where
$\mathbf{P}(u)=P(D)\left(u^{2}+\frac{1}{2}(\partial_{x}u)^{2}\right)\quad\text{with}\quad P(D)=-\partial_{x}\Lambda^{-2}.$ (1.3)
The CH equation (1.1) was firstly proposed in the context of hereditary symmetries studied in [16] and then was derived explicitly as a water wave equation by Camassa–Holm [2]. Many aspects of the mathematical beauty of the CH equation have been exposed over the last two decades. Particularly, (CH) is completely integrable [2, 5] with a bi-Hamiltonian structure [4, 16] and infinitely many conservation laws [2, 16]. Also, it admits exact peaked soliton solutions (peakons) of the form $u(x,t)=ce^{-|x-ct|}\;(c>0)$, which are orbitally stable [12]. Another remarkable feature of the CH equation is the wave breaking phenomena: the solution remains bounded while its slope becomes unbounded in finite time [3, 8, 9]. It is worth mentioning that the peaked solitons present the characteristic for the travelling water waves of greatest height and largest amplitude and arise as solutions to the free-boundary problem for incompressible Euler equations over a flat bed, see Refs. [6, 10, 11, 39] for the details.
Due to these interesting and remarkable features, the CH equation has attracted much attention as a class of integrable shallow water wave equations in recent twenty years. Its systematic mathematical study was initiated in a series of papers by Constantin and Escher, see [7, 8, 9, 10, 11]. After the CH equation was derived physically in the context of water waves, there are a large amount of literatures devoted to studying the well-posedness of the Cauchy problem (1.1) (see Molinet’s survey [37]). Li and Olver [36] proved that the Cauchy problem (1.1) is locally well-posed with the initial data $u_{0}\in H^{s}(\mathbb{R})$ with $s>3/2$ (see also [38]). Danchin [13, 14] proved the local existence and uniqueness of strong solutions to (1.1) with initial data in $B^{s}_{p,r}$ if $(p,r)\in[1,\infty]\times[1,\infty),s>\max\big{\{}1+1/p,3/2\big{\}}$ and $B^{3/2}_{2,1}$. Meanwhile, he [13] only obtained the continuity of the solution map of (1.1) with respect to the initial data in the space $\mathcal{C}([0,T];B^{s^{\prime}}_{p,r})$ with any $s^{\prime}<s$. Li-Yin [32] proved the continuity of the solution map of (1.1) with respect to the initial data in the space $\mathcal{C}([0,T];B^{s}_{p,r})$ with $r<\infty$. In particular, they [32] proved that the solution map of (1.1) is weak continuous with respect to initial data $u_{0}\in B^{s}_{p,\infty}$. For the endpoints, Danchin [14] obtained that the data-to-solution map is not continuous by using peakon solution, which implies the ill-posedness of (1.1) in $B^{3/2}_{2,\infty}$. Himonas-Misiołek [23] obtained the first result on the non-uniform dependence for (1.1) in $H^{s}(\mathbb{T})$ with $s\geq 2$ using explicitly constructed travelling wave solutions, which was sharpened to $s>\frac{3}{2}$ by Himonas-Kenig [21] on the real-line and Himonas-Kenig-Misiołek [22] on the circle. We should mention that, non-uniform continuity of the CH solution map in $H^{1}(\mathbb{R}\;\text{or}\;\mathbb{T})$ was established by Himonas-Misiołek-Ponce [24] by using traveling wave solutions. In our recent papers [31, 30], we proved the non-uniform dependence on initial data for (1.1) under both the framework of Besov spaces $B^{s}_{p,r}$ for $s>\max\big{\{}1+1/p,3/2\big{\}}$ with $(p,r)\in[1,\infty]\times[1,\infty)$ and $B^{3/2}_{2,1}$. Guo-Liu-Molinet-Yin [17] showed the ill-posedness of (1.1) in $B_{p,r}^{1+1/p}(\mathbb{R}\;\text{or}\;\mathbb{T})$ with $(p,r)\in[1,\infty]\times(1,\infty]$ (especially in $H^{3/2}$) by proving the norm inflation. Very recently, Guo-Ye-Yin [18] obtained the ill-posedness for the CH equation in $B^{1}_{\infty,1}(\mathbb{R})$ by proving the norm inflation. Ye-Yin-Guo [40] obtained the local well-posedness for the Camassa–Holm type equation in Besov spaces $s>1+\frac{1}{p},(p,r)\in[1,+\infty]\times[1,+\infty)$ or $s=1+\frac{1}{p},\ (p,r)\in[1,+\infty)\times\{1\}$. In our recent papers [33, 34], we established the ill-posedness for (1.1) in $B^{s}_{p,\infty}(\mathbb{R})$ with $s>\max\big{\{}1+1/p,3/2\big{\}}$ with $p\in[1,\infty]$ by proving the solution map to the CH equation starting from $u_{0}$ is discontinuous at $t=0$ in the metric of $B^{s}_{p,\infty}(\mathbb{R})$.
Assume that $u_{0}\in B^{s}_{p,r}$ with $s>1+\frac{1}{p},(p,r)\in[1,+\infty]\times[1,+\infty)$ or $s=1+\frac{1}{p},\ (p,r)\in[1,+\infty)\times\{1\}$, it is known that there exists a solution $u\in\mathcal{C}([0,T];B^{s}_{p,r})$ for the Camassa-Holm solution. Naturally, we may wonder whether or not the solution $u$ can belong to $\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with some $\alpha\in(0,1)$. Therefore, we are interested in the following question:
$u_{0}\in B^{s}_{p,r}\;\Rightarrow\;\exists 1\;u\in\mathcal{C}([0,T];B^{s}_{p,r})\;\overset{?}{\Rightarrow}\;\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})\quad\text{with}\;\alpha\in(0,1).$
If the initial data $u_{0}$ have more regularity such that $u_{0}\in B^{s^{\prime}}_{p,r}$ for some $s^{\prime}>s$, by the interpolation argument, we can deduce that $u\in\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with $\alpha=s^{\prime}-s$. In this paper, we will show that there exits initial data $u_{0}\in B^{s}_{p,r}$ such that the corresponding solution of the Camassa-Holm solution can not belong to $\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with any $\alpha\in(0,1)$. Now let us state our main result of this paper.
Theorem 1.1.
Assume that $(s,p,r)$ satisfies that
[…]
For any $\alpha\in(0,1)$, there exits $u_{0}\in B^{s}_{p,r}(\mathbb{R})$ such that the data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})\in\mathcal{C}([0,T];B^{s}_{p,r})$ of the Cauchy problem (1.2) satisfies
$\limsup_{t\to 0^{+}}\frac{\|\mathbf{S}_{t}(u_{0})-u_{0}\|_{B^{s}_{p,r}}}{t^{\alpha}}=+\infty.$
Remark 1.1.
Theorem 1.1 also holds for the following b-family equation (see [25, 26, 15, 27, 28] etc.):
$\begin{cases}\partial_{t}u+uu_{x}=-\partial_{x}(1-\partial^{2}_{x})^{-1}\big{(}\frac{b}{2}u^{2}+\frac{3-b}{2}u^{2}_{x}\big{)},&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R},\\ u(0,x)=u_{0}(x),&\quad x\in\mathbb{R}.\end{cases}$ (1.5)
It should be mentioned that the Camassa-Holm equation corresponds to $b=2$ and Degasperis-Procesi equation corresponds to $b=3$ [35].
Remark 1.2.
We summary the local well-posedness/ill-posdeness results of the Camassa-Holm equation in the Besov spaces. This can be seen clearly from the Table below.
References Range Results
[13, 14, 32] $s>\max\big{\{}1+\frac{1}{p},\frac{3}{2}\big{\}},(p,r)\in[1,\infty]\times[1,\infty)$ and $s=\frac{3}{2},p=2,r=1$ LWP
[17] $s=1+\frac{1}{p},(p,r)\in[1,\infty]\times(1,\infty]$ Norm inflation
[18] $s=1,p=\infty,r=1$ Norm inflation
[33, 34] $s>\max\big{\{}1+\frac{1}{p},\frac{3}{2}\big{\}},p\in[1,\infty],r=\infty$ Discontinuos
[40] $s>1+\frac{1}{p},p\in[1,+\infty],r\in[1,+\infty)$ and $s=1+\frac{1}{p},p\in[1,+\infty),r=1$ LWP
$s>1+\frac{1}{p},p\in[1,+\infty],r\in[1,+\infty)$ and $s=1+\frac{1}{p},p\in[1,+\infty),r=1$ Theorem 1.1
Table 1: Well/Ill-posedness of (1.2) in $B^{s}_{p,r}$
The Cauchy problem for the Novikov equation reads as (see [19, 20, 29] etc.)
$\begin{cases}v_{t}+v^{2}v_{x}=-(1-\partial^{2}_{x})^{-1}\Big{(}\frac{1}{2}v_{x}^{3}+\partial_{x}\big{(}\frac{3}{2}vv^{2}_{x}+v^{3}\big{)}\Big{)}=:\mathbf{Q}(v),\\ v(0,x)=v_{0}.\end{cases}$ (1.6)
Then, we have the following
Theorem 1.2.
Assume that $(s,p,r)$ satisfies that (1.4). For any $\alpha\in(0,1)$, there exits $v_{0}\in B^{s}_{p,r}(\mathbb{R})$ such that the data-to-solution map $v_{0}\mapsto\mathbf{S}_{t}(v_{0})\in\mathcal{C}([0,T];B^{s}_{p,r})$ of the Cauchy problem (1.6) satisfies
$\limsup_{t\to 0^{+}}\frac{\|\mathbf{S}_{t}(v_{0})-v_{0}\|_{B^{s}_{p,r}}}{t^{\alpha}}=+\infty.$
Remark 1.3.
Following out method, Theorem 1.1 and Theorem 1.2 also hold for the Torus case. We omit the details and leave to the interesting readers.
2 Preliminaries
Notation  The notation $A\lesssim B$ (resp., $A\gtrsim B$) means that there exists a harmless positive constant $c$ such that $A\leq cB$ (resp., $A\geq cB$). Given a Banach space $X$, we denote its norm by $\|\cdot\|_{X}$. For $I\subset\mathbb{R}$, we denote by $\mathcal{C}(I;X)$ the set of continuous functions on $I$ with values in $X$. Sometimes we will denote $L^{p}(0,T;X)$ by $L_{T}^{p}X$.
Let us recall that for all $f\in\mathcal{S}^{\prime}$, the Fourier transform $\widehat{f}$, is defined by
$(\mathcal{F}f)(\xi)=\widehat{f}(\xi)=\int_{\mathbb{R}}e^{-\mathrm{i}x\xi}f(x)\mathrm{d}x\quad\text{for any}\;\xi\in\mathbb{R}.$
Next, we will recall some facts about the Littlewood-Paley decomposition and the nonhomogeneous Besov spaces (see [1] for more details). Choose a radial, non-negative, smooth function $\vartheta:\mathbb{R}\mapsto[0,1]$ such that ${\rm{supp}}\,\vartheta\subset B(0,4/3)$ and $\vartheta(\xi)\equiv 1$ for $|\xi|\leq 3/4$. Setting $\varphi(\xi):=\vartheta(\xi/2)-\vartheta(\xi)$, then we deduce that $\varphi$ has the following properties
  • ${\rm{supp}}\;\varphi\subset\left\{\xi\in\mathbb{R}:3/4\leq|\xi|\leq 8/3\right\}$;
  • $\varphi(\xi)\equiv 1$ for $24/3\leq|\xi|\leq 3/$ 2;
  • $\vartheta(\xi)+\sum_{j\geq 0}\varphi(2^{-j}\xi)=1$ for any $\xi\in\mathbb{R}$.
Definition 2.1 (see [1]).
For every $u\in\mathcal{S^{\prime}}(\mathbb{R})$, the Littlewood-Paley dyadic blocks ${\Delta}_{j}$ are defined as follows
$\Delta_{j}u=$ $0,$ if $j\leq-2$;
$\Delta_{j}u=$ $\chi(D)u=\mathcal{F}^{-1}(\chi\mathcal{F}u),$ if $j=-1$;
$\Delta_{j}u=$ $\varphi(2^{-j}D)u=\mathcal{F}^{-1}\big{(}\varphi(2^{-j}\cdot)\mathcal{F}u\big{)},$ if $j\geq 0$.
The inhomogeneous low-frequency cut-off operator $S_{j}$ is defined by
$S_{j}u=\sum_{q=-1}^{j-1}{\Delta}_{q}u.$
Definition 2.2 (see [1]).
Let $s\in\mathbb{R}$ and $(p,r)\in[1,\infty]^{2}$. The nonhomogeneous Besov space $B^{s}_{p,r}(\mathbb{R})$ is defined by
$B^{s}_{p,r}(\mathbb{R}):=\Big{\{}f\in\mathcal{S}^{\prime}(\mathbb{R}):\;\|f\|_{B^{s}_{p,r}(\mathbb{R})}<\infty\Big{\}},$
where
$\|f\|_{B^{s}_{p,r}(\mathbb{R})}=$ $\left(\sum_{j\geq-1}2^{sjr}\|\Delta_{j}f\|^{r}_{L^{p}(\mathbb{R})}\right)^{\frac{1}{r}},$ if $1\leq r<\infty$,
$\|f\|_{B^{s}_{p,r}(\mathbb{R})}=$ $\sup_{j\geq-1}2^{sj}\|\Delta_{j}f\|_{L^{p}(\mathbb{R})},$ if $r=\infty$.
We shall use Bony’s decomposition [1] in the nonhomogeneous context throughout this paper
$uv={T}_{u}v+{T}_{v}u+{R}(u,v)\quad\text{with}$
${T}_{u}v=\sum_{j\geq-1}{S}_{j-1}u{\Delta}_{j}v\quad\mbox{and}\quad{R}(u,v)=\sum_{|j-k|\leq 1}{\Delta}_{j}u{\Delta}_{k}v.$
Lemma 2.1 (see [1]).
Let $\left(s,t,s_{1},s_{2}\right)\in\mathbb{R}^{4}$ and $\left(p,p_{1},p_{2},r,r_{1},r_{2}\right)\in[1,\infty]^{6}$. Assume that
$\frac{1}{p}=\frac{1}{p_{1}}+\frac{1}{p_{2}}\leq 1,\quad\frac{1}{r}=\frac{1}{r_{1}}+\frac{1}{r_{2}}\leq 1,\quad s_{1}+s_{2}>0,\quad t<0.$
There exists a constant $C$ such that
$\left\|T_{u}v\right\|_{B_{p,r}^{s}\left(\mathbb{R}^{d}\right)}\leq C^{|s|+1}\|u\|_{L^{\infty}\left(\mathbb{R}^{d}\right)}\|v\|_{B_{p,r}^{s}\left(\mathbb{R}^{d}\right)},$
$\left\|T_{u}v\right\|_{B_{p,r}^{s+t}\left(\mathbb{R}^{d}\right)}\leq\frac{C^{|s+t|+1}}{-t}\|u\|_{B_{\infty,r_{1}}^{t}\left(\mathbb{R}^{d}\right)}\|v\|_{B_{p,r_{2}}^{s}\left(\mathbb{R}^{d}\right)},$
$\|R(u,v)\|_{B_{p,r}^{s_{1}+s_{2}}\left(\mathbb{R}^{d}\right)}\leq\frac{C^{\left|s_{1}+s_{2}\right|+1}}{s_{1}+s_{2}}\|u\|_{B_{p_{1},r_{1}}^{s_{1}}\left(\mathbb{R}^{d}\right)}\|v\|_{B_{p_{2},r_{2}}^{s_{2}}\left(\mathbb{R}^{d}\right)}.$
Finally, we give some important properties which will be also often used throughout the paper.
Lemma 2.2 (see [1]).
Let $(p,r)\in[1,\infty]^{2}$ and $s>\max\big{\{}1+\frac{1}{p},\frac{3}{2}\big{\}}$. Then we have
$\|uv\|_{B^{s-2}_{p,r}(\mathbb{R})}\leq C\|u\|_{B^{s-2}_{p,r}(\mathbb{R})}\|v\|_{B^{s-1}_{p,r}(\mathbb{R})}.$
Lemma 2.3 (see [1]).
For $(p,r)\in[1,\infty]^{2}$, $B^{s-1}_{p,r}(\mathbb{R})$ with $s>1+\frac{1}{p}$ or $s=1+\frac{1}{p},\ r=1$ is an algebra. Moreover, for any $u,v\in B^{s-1}_{p,r}(\mathbb{R})$ with $s>1+\frac{1}{p}$ or $s=1+\frac{1}{p},\ r=1$ , we have
$\|uv\|_{B^{s-1}_{p,r}(\mathbb{R})}\leq C\|u\|_{B^{s-1}_{p,r}(\mathbb{R})}\|v\|_{B^{s-1}_{p,r}(\mathbb{R})}.$
Remark 2.1.
Let $(p,r)\in[1,\infty]^{2}$ and $s>\max\big{\{}1+\frac{1}{p},\frac{3}{2}\big{\}}$, using Lemmas 2.2-2.3, we have
$\|\mathbf{P}(u)-\mathbf{P}(v)\|_{B_{p,r}^{s-1}}$ $\lesssim\|u-v\|_{B_{p,r}^{s-1}}\big{(}\|u\|_{B_{p,r}^{s}}+\|v\|_{B_{p,r}^{s}}\big{)}.$ (2.7)
Lemma 2.4 (see [1]).
For $(p,r)\in[1,\infty]^{2}$, $s_{1}\neq s_{2}$ and $\theta\in(0,1)$, the following inequality holds
$\|u\|_{{B}_{p,r}^{\theta s_{1}+(1-\theta)s_{2}}}\leq\|u\|^{\theta}_{{B}_{p,r}^{s_{1}}}\|u\|^{1-\theta}_{{B}_{p,r}^{s_{2}}}.$
Lemma 2.5 (see [1]).
For $1\leq p\leq\infty$ and $s>0$. There exists a constant $C$, depending continuously on $p$ and $s$, we have
$\left\|2^{js}\left\|[\Delta_{j},v]\partial_{x}f\right\|_{L^{p}}\right\|_{\ell^{\infty}}\leq C\big{(}\|\partial_{x}v\|_{L^{\infty}}\|f\|_{B_{p,\infty}^{s}}+\|\partial_{x}f\|_{L^{\infty}}\|\partial_{x}v\|_{B_{p,\infty}^{s-1}}\big{)},$
where we denote the standard commutator $[\Delta_{j},v]\partial_{x}f=\Delta_{j}(v\partial_{x}f)-v\Delta_{j}\partial_{x}f$.
3 Proof of Main Theorems
3.1 Construction of Initial Data
We need to introduce smooth, radial cut-off functions to localize the frequency region. Precisely, let $\widehat{\phi}\in\mathcal{C}^{\infty}_{0}(\mathbb{R})$ be an even, real-valued and non-negative function on $\mathbb{R}$ and satisfy
$\widehat{\phi}(\xi)=$ $1,$ if $|\xi|\leq\frac{1}{4}$,
$\widehat{\phi}(\xi)=$ $0,$ if $|\xi|\geq\frac{1}{2}$.
Obviously, $\phi(0)>0$ and for any $p\in[1,\infty]$, there exists two positive constants $c_{1}$ and $c_{2}$ such that
$c_{1}\leq\|\phi\|_{L^{p}(\mathbb{R})}\leq c_{2}.$
Motivated by [31], we define the function $f_{n}(x)$ by
$f_{n}(x)=\phi(x)\cos\left(\frac{17}{12}2^{n}x\right)\quad\text{with}\quad n\gg 1.$
Due to the fact $\varphi(\xi)\equiv 1$ for $\frac{4}{3}\leq|\xi|\leq\frac{3}{2}$, namely,
$\varphi(2^{-j}\xi)\equiv 1\quad\text{in}\;\left\{\xi\in\mathbb{R}^{d}:\ \frac{4}{3}2^{j}\leq|\xi|\leq\frac{3}{2}2^{j}\right\},$
we have
${\Delta_{j}(f_{n})=\mathcal{F}^{-1}\left(\varphi(2^{-j}\cdot)\widehat{f}_{n}\right)=}\begin{cases}f_{n},&\text{if}\;j=n,\\ 0,&\text{if}\;j\neq n.\end{cases}$ (3.8)
Lemma 3.1.
Assume that $(s,p,r)$ satisfies (1.4). Define the initial data $u_{0}(x)$ as
$u_{0}(x):=\sum\limits^{\infty}_{n=3}n^{-2}2^{-ns}\phi(x)\cos\left(\frac{17}{12}2^{n}x\right).$ (3.9)
Then there exists some sufficiently large $n\in\mathbb{Z}^{+}$ and some sufficiently enough $\delta>0$ such that
$\|u_{0}\|_{B^{s}_{p,r}}\leq C,$
$\|u_{0}^{k}\partial_{x}\Delta_{n}u_{0}\|_{L^{p}}\geq cn^{-2}2^{n(1-s)},\quad k\in\{1,2\}$
where $C$ and $c$ are some positive constants.
Proof.  Using (3.8) yields
$\Delta_{n}u_{0}(x)$ $=n^{-2}2^{-ns}\phi(x)\cos\left(\frac{17}{12}2^{n}x\right).$ (3.10)
By the definition of ${B}_{p,r}^{s}$ and , we deduce that
$\|u_{0}\|_{{B}_{p,r}^{s}(\mathbb{R})}$ $=\left\|2^{js}\|\Delta_{j}u_{0}\|_{L^{p}(\mathbb{R})}\right\|_{\ell^{r}(j\geq 1)}\leq\left\|\frac{1}{j^{2}}\right\|_{\ell^{r}(j\geq 1)}\|\phi\|_{L^{p}(\mathbb{R})}\leq C\|\phi\|_{L^{p}(\mathbb{R})}.$
From (3.10), we have
$n^{2}u_{0}^{k}\partial_{x}\Delta_{n}u_{0}$ $=2^{-ns}u_{0}^{k}(x)\phi^{\prime}(x)\cos\left(\frac{17}{12}2^{n}x\right)-\frac{17}{12}2^{n}2^{-ns}u_{0}^{k}(x)\phi(x)\sin\left(\frac{17}{12}2^{n}x\right).$
Since $u_{0}^{k}(x)$ is a real-valued and continuous function on $\mathbb{R}$, then there exists some $\delta>0$ such that
$|u_{0}^{k}(x)|\geq\frac{1}{2}|u_{0}^{k}(0)|=\frac{1}{2}\left(\phi(0)\sum\limits^{\infty}_{n=3}n^{-2}2^{-ns}\right)^{k}=:c_{0}\quad\text{ for any }x\in B_{\delta}(0).$ (3.11)
Thus we have from (3.11)
$n^{2}\|u_{0}^{k}\partial_{x}\Delta_{n}u_{0}\|_{L^{p}}$ $\geq c_{0}2^{n}2^{-ns}\left\|\phi(x)\sin\left(\frac{17}{12}2^{n}x\right)\right\|_{L^{p}(B_{\delta}(0))}-C2^{-ns}\left\|\phi^{\prime}(x)\phi^{k}(x)\cos\left(\frac{17}{12}2^{n}x\right)\right\|_{L^{p}}$
$\geq(c2^{n}-C)2^{-ns}.$
We choose $n$ large enough such that $C<\frac{c}{2}2^{n}$ and then finish the proof of Lemma 3.1.
3.2 Error Estimates
Proposition 3.1.
Assume that $u_{0}$ satisfies (3.9). Under the assumptions of Theorem 1.1, we have
$\|\mathbf{S}_{t}\left(u_{0}\right)-u_{0}\|_{B^{s-1}_{p,r}}\lesssim t.$ (3.12)
Furthermore, there holds
  • for $s>\max\left\{1+\frac{1}{p},\frac{3}{2}\right\}$ and $(p,r)\in[1,\infty]\times[1,\infty),$ we have
    $\|\mathbf{w}\|_{B^{s-2}_{p,r}}\lesssim t^{2};$
    here and in what follows we denote
[…]
  • $1+\frac{1}{p}<s\leq\frac{3}{2}$ and $(p,r)\in[1,\infty]\times[1,\infty)$ or $s=1+\frac{1}{p}$ and $(p,r)\in[1,\infty)\times\{1\}$, we have
    $\|\mathbf{w}\|_{B^{0}_{p,r}}\lesssim t^{s}.$
Proof.  For simplicity, we denote $u(t):=\mathbf{S}_{t}(u_{0})$ here and in what follows. Notice that $(s,p,r)$ satisfies (1.4), then using the local well-posedness result, we know that there exists a positive time $T$ such that $u(t)\in\mathcal{C}([0,T];B_{p,r}^{s})$. Furthermore, it holds that
$\|u(t)\|_{L^{\infty}_{T}B^{s}_{p,r}}\leq C\|u_{0}\|_{B^{s}_{p,r}}\leq C.$
Noticing that $u_{0}\in B^{s}_{\infty,r}$, we can deduce that $u(t)\in\mathcal{C}([0,T];B_{\infty,r}^{s})$ and also hold
$\|u(t)\|_{L^{\infty}_{T}B^{s}_{\infty,r}}\leq C\|u_{0}\|_{B^{s}_{\infty,r}}\leq C.$
Using the Newton-Leibniz formula and Remark 2.1 with $v=0$, we obtain from (1.2) that
$\|u(t)-u_{0}\|_{B^{s-1}_{p,r}}$ $\leq\int^{t}_{0}\|\partial_{\tau}u\|_{B^{s-1}_{p,r}}\mathrm{d}\tau$
$\leq\int^{t}_{0}\|\mathbf{P}(u)\|_{B^{s-1}_{p,r}}\mathrm{d}\tau+\int^{t}_{0}\|u\partial_{x}u\|_{B^{s-1}_{p,r}}\mathrm{d}\tau$
$\lesssim t\|u\|^{2}_{L_{t}^{\infty}B^{s}_{p,r}}$
$\lesssim t\|u_{0}\|^{2}_{B^{s}_{p,r}}$
t.less-than-or-similar-$\lesssim t.$ (3.13)
Case 1: $s>\max\left\{1+\frac{1}{p},\frac{3}{2}\right\}$ and $(p,r)\in[1,\infty]\times[1,\infty)$. By the Newton-Leibniz formula and Lemmas 2.2-2.3, we obtain from (3.2) that
$\|\mathbf{w}\|_{B^{s-2}_{p,r}}$ $\leq\int^{t}_{0}\|\partial_{\tau}u-\mathbf{v}_{0}\|_{B^{s-2}_{p,r}}\mathrm{d}\tau$
$\lesssim\int^{t}_{0}\|\mathbf{P}(u)-\mathbf{P}(u_{0})\|_{B^{s-2}_{p,r}}\mathrm{d}\tau+\int^{t}_{0}\|u\partial_{x}u-u_{0}\partial_{x}u_{0}\|_{B^{s-2}_{p,r}}\mathrm{d}\tau$
$\lesssim\int^{t}_{0}\|\partial_{x}(u-u_{0})\partial_{x}(u+u_{0})\|_{B^{s-2}_{p,r}}\mathrm{d}\tau+\int^{t}_{0}\|(u-u_{0})(u+u_{0})\|_{B^{s-1}_{p,r}}\mathrm{d}\tau$
$\lesssim\int^{t}_{0}\|\partial_{x}(u-u_{0})\|_{B^{s-2}_{p,r}}\|\partial_{x}(u+u_{0})\|_{B^{s-1}_{p,r}}\mathrm{d}\tau+\int^{t}_{0}\|u-u_{0}\|_{B^{s-1}_{p,r}}\|u+u_{0}\|_{B^{s-1}_{p,r}}\mathrm{d}\tau$
$\lesssim t^{2}.$ (3.14)
Case 2: $1+\frac{1}{p}<s\leq\frac{3}{2}$ and $(p,r)\in[1,\infty]\times[1,\infty)$ or $s=1+\frac{1}{p}$ and $(p,r)\in[1,\infty)\times\{1\}$.
Then using the embedding $B^{s-2}_{p,r}\hookrightarrow B^{-1}_{p,r}$, we have
$\|\mathbf{w}\|_{B^{0}_{p,r}}$ $\leq\int^{t}_{0}\|\partial_{\tau}u-\mathbf{v}_{0}\|_{B^{0}_{p,r}}\mathrm{d}\tau$
$\lesssim\int^{t}_{0}\|\mathbf{P}(u)-\mathbf{P}(u_{0})\|_{B^{0}_{p,r}}\mathrm{d}\tau+\int^{t}_{0}\|u\partial_{x}u-u_{0}\partial_{x}u_{0}\|_{B^{0}_{p,r}}\mathrm{d}\tau$
$\lesssim\int^{t}_{0}\|\partial_{x}(u-u_{0})\partial_{x}(u+u_{0})\|_{B^{-1}_{p,r}}\mathrm{d}\tau+\int^{t}_{0}\|(u-u_{0})(u+u_{0})\|_{B^{1}_{p,r}}\mathrm{d}\tau$
$\lesssim\int^{t}_{0}\underbrace{\|\partial_{x}(u-u_{0})\partial_{x}(u+u_{0})\|_{B^{s-2}_{p,r}}}_{=:\ I_{1}}\mathrm{d}\tau+\int^{t}_{0}\underbrace{\|u-u_{0}\|_{B^{1}_{p,r}}\|u+u_{0}\|_{B^{1}_{p,r}}}_{=:\ I_{2}}\mathrm{d}\tau.$ (3.15)
To estimate the term $I_{1}$, by Bony’s decomposition, one has
$\partial_{x}(u-u_{0})\partial_{x}(u+u_{0})=T_{\partial_{x}(u-u_{0})}\partial_{x}(u+u_{0})+T_{\partial_{x}(u+u_{0})}\partial_{x}(u-u_{0})+R\big{(}\partial_{x}(u+u_{0}),\partial_{x}(u-u_{0})\big{)}.$
Using Lemma 3.1 yields that
$\|T_{\partial_{x}(u-u_{0})}\partial_{x}(u+u_{0})\|_{B^{s-2}_{p,r}}$ $\lesssim\|\partial_{x}(u-u_{0})\|_{B^{-1}_{\infty,\infty}}\|\partial_{x}(u+u_{0})\|_{B^{s-1}_{p,r}}$
$\lesssim\|u-u_{0}\|_{B^{1}_{p,r}}\|u+u_{0}\|_{B^{s}_{p,r}}$
$\lesssim t^{s-1},$
$\|T_{\partial_{x}(u+u_{0})}\partial_{x}(u-u_{0})\|_{B^{s-2}_{p,r}}$ $\lesssim\|\partial_{x}(u+u_{0})\|_{L^{\infty}}\|\partial_{x}(u-u_{0})\|_{B^{s-2}_{p,r}}$
$\lesssim\|u+u_{0}\|_{B^{s-1}_{p,r}}\|u-u_{0}\|_{B^{s-1}_{p,r}}$
t,less-than-or-similar-$\lesssim t,$
$\|R\big{(}\partial_{x}(u+u_{0}),\partial_{x}(u-u_{0})\big{)}\|_{B^{s-1}_{p,r}}$ $\lesssim\|\partial_{x}(u+u_{0})\|_{B^{s-1}_{\infty,\infty}}\|\partial_{x}(u-u_{0})\|_{B^{0}_{p,r}}$
$\lesssim\|u+u_{0}\|_{B^{s}_{\infty,r}}\|u-u_{0}\|_{B^{1}_{p,r}}$
$\lesssim t^{s-1},$
where we have used the fact
$\|u(t)\|_{B^{s}_{\infty,r}}\lesssim\|u_{0}\|_{B^{s}_{\infty,r}}\lesssim 1$
and the interpolation argument
$\|u-u_{0}\|_{B^{1}_{p,r}}\lesssim\|u(t)-u_{0}\|^{s-1}_{B^{s-1}_{p,r}}\|u(t)-u_{0}\|^{2-s}_{B^{s}_{p,r}}\lesssim t^{s-1}.$
Combining the above, we obtain
$I_{1}\lesssim t^{s-1}.$ (3.16)
To estimate the term $I_{2}$, by the interpolation argument, one has
$I_{2}\lesssim\|u-u_{0}\|_{B^{1}_{p,r}}\lesssim t^{s-1}.$ (3.17)
Inserting (3.16) and (3.17) into (3.2) yields the desired result. Thus, we finish the proof of Proposition 3.1.
3.3 Proof of Theorem 1.1
Now we present the proof of Theorem 1.1. Notice that
$\mathbf{S}_{t}(u_{0})-u_{0}=t\mathbf{\widetilde{u}}_{0}+\mathbf{w}\quad\text{and}\quad\mathbf{\widetilde{u}}_{0}=\mathbf{P}(u_{0})-u_{0}\partial_{x}u_{0}.$
Case 1: $s>\max\left\{1+\frac{1}{p},\frac{3}{2}\right\}$ and $(p,r)\in[1,\infty]\times[1,\infty)$.
By the triangle inequality and Propositions 3.1, we deduce that
$\|\mathbf{S}_{t}(u_{0})-u_{0}\|_{B^{s}_{p,r}}$ $\geq 2^{{ns}}\big{\|}\Delta_{n}\big{(}\mathbf{S}_{t}(u_{0})-u_{0}\big{)}\big{\|}_{L^{p}}$
$=2^{{ns}}\big{\|}\Delta_{n}\big{(}t\mathbf{\widetilde{u}}_{0}+\mathbf{w}\big{)}\big{\|}_{L^{p}}$
$\geq t2^{{ns}}\|\Delta_{n}\mathbf{\widetilde{u}}_{0}\|_{L^{p}}-2^{{2n}}2^{{n(s-2)}}\big{\|}\Delta_{n}\mathbf{w}\big{\|}_{L^{p}}$
$\geq t2^{{n}s}\|\Delta_{n}\big{(}u_{0}\partial_{x}u_{0}\big{)}\|_{L^{p}}-t2^{{n}s}\|\Delta_{n}\big{(}\mathbf{P}(u_{0})\big{)}\|_{L^{p}}-C2^{2{n}}\|\mathbf{w}\|_{B^{s-2}_{p,\infty}}$
$\geq t2^{{n}s}\|u_{0}\partial_{x}\Delta_{n}u_{0}\|_{L^{p}}-t2^{{n}s}\|[\Delta_{n},u_{0}]\partial_{x}u_{0}\|_{L^{p}}$
$\quad-Ct\|\mathbf{P}(u_{0})\|_{B^{s}_{p,\infty}}-C2^{2{n}}t^{2}$
$\geq t2^{{n}s}\|u_{0}\partial_{x}\Delta_{n}u_{0}\|_{L^{p}}-Ct\big{\|}2^{{n}s}\|[\Delta_{n},u_{0}]\partial_{x}u_{0}\|_{L^{p}}\big{\|}_{\ell^{\infty}}$
$\quad-t\|\mathbf{P}(u_{0})\|_{B^{s}_{p,\infty}}-C2^{2{n}}t^{2}.$ (3.18)
By Lemmas 2.2-2.3, one has
$\|\mathbf{P}(u_{0})\|_{B^{s}_{p,\infty}}\lesssim\|(\partial_{x}u_{0})^{2}+u_{0}^{2}\|_{B^{s-1}_{p,r}}\lesssim 1$
and by Lemma 2.5
$\big{\|}2^{{n}s}\|[\Delta_{n},u_{0}]\partial_{x}u_{0}\|_{L^{p}}\big{\|}_{\ell^{\infty}}\lesssim\|\partial_{x}u_{0}\|_{L^{\infty}}\|u_{0}\|_{B^{s}_{p,\infty}}+\|\partial_{x}u_{0}\|_{L^{\infty}}\|\partial_{x}u_{0}\|_{B^{s-1}_{p,\infty}}\lesssim 1.$
Gathering all the above estimates and Lemma 3.1 together with (3.3), we obtain
$\|\mathbf{S}_{t}(u_{0})-u_{0}\|_{B^{s}_{p,r}}\geq ctn^{-2}2^{n}-Ct-C2^{2{n}}t^{2},$
which implies
$t^{-\alpha}\|\mathbf{S}_{t}(u_{0})-u_{0}\|_{B^{s}_{p,r}}\geq ct^{1-\alpha}n^{-2}2^{n}-Ct^{1-\alpha}-C2^{2{n}}t^{2-\alpha},$
Thus, picking $t^{1-\alpha}_{n}=n^{3}2^{-n}$ with large $n$, we have
$t^{-\alpha}_{n}\|\mathbf{S}_{t_{n}}(u_{0})-u_{0}\|_{B^{s}_{p,r}}$ $\geq cn-Cn^{3}2^{-n}-Cn^{6}t_{n}^{\alpha}\geq\tilde{c}n.$
Case 2: $1+\frac{1}{p}<s\leq\frac{3}{2}$ and $(p,r)\in[1,\infty]\times[1,\infty)$ or $s=1+\frac{1}{p}$ and $(p,r)\in[1,\infty)\times\{1\}$.
By the triangle inequality and Propositions 3.1, we deduce that
$\|\mathbf{S}_{t}(u_{0})-u_{0}\|_{B^{s}_{p,r}}$ $\geq 2^{{ns}}\big{\|}\Delta_{n}\big{(}\mathbf{S}_{t}(u_{0})-u_{0}\big{)}\big{\|}_{L^{p}}$
$=2^{{ns}}\big{\|}\Delta_{n}\big{(}t\mathbf{v}_{0}+\mathbf{w}\big{)}\big{\|}_{L^{p}}$
$\geq t2^{{ns}}\|\Delta_{n}\mathbf{v}_{0}\|_{L^{p}}-2^{{ns}}\|\mathbf{w}\|_{B^{0}_{p,r}}$
$\geq t2^{{n}s}\|\Delta_{n}\big{(}u_{0}\partial_{x}u_{0}\big{)}\|_{L^{p}}-t2^{{n}s}\|\Delta_{n}\big{(}\mathbf{P}(u_{0})\big{)}\|_{L^{p}}-C(2^{{n}}t)^{s}$
$\geq t2^{{n}s}\|u_{0}\partial_{x}\Delta_{n}u_{0}\|_{L^{p}}-t2^{{n}s}\|[\Delta_{n},u_{0}]\partial_{x}u_{0}\|_{L^{p}}$
$\quad-t\|\mathbf{P}(u_{0})\|_{B^{s}_{p,r}}-C(2^{{n}}t)^{s}$
$\geq t2^{{n}s}\|u_{0}\partial_{x}\Delta_{n}u_{0}\|_{L^{p}}-Ct\big{\|}2^{{n}s}\|[\Delta_{n},u_{0}]\partial_{x}u_{0}\|_{L^{p}}\big{\|}_{\ell^{\infty}}$
$\quad-t\|\mathbf{P}(u_{0})\|_{B^{s}_{p,\infty}}-C(2^{{n}}t)^{s},$ (3.19)
where we have used
$\|\mathbf{P}(u_{0})\|_{B^{s}_{p,r}}\lesssim\|(\partial_{x}u_{0})^{2}+u_{0}^{2}\|_{B^{s}_{p,r}}\lesssim 1$
and
$\big{\|}2^{{n}s}\|[\Delta_{n},u_{0}]\partial_{x}u_{0}\|_{L^{p}}\big{\|}_{\ell^{\infty}}\lesssim\|\partial_{x}u_{0}\|_{L^{\infty}}\|u_{0}\|_{B^{s}_{p,\infty}}+\|\partial_{x}u_{0}\|_{L^{\infty}}\|\partial_{x}u_{0}\|_{B^{s-1}_{p,\infty}}\lesssim 1.$
Gathering Lemma 3.1 together with (3.3), we obtain
$\|\mathbf{S}_{t}(u_{0})-u_{0}\|_{B^{s}_{p,r}}\geq ctn^{-2}2^{{n}}-Ct-C(2^{{n}}t)^{s},$
which implies
$t^{-\alpha}\|\mathbf{S}_{t}(u_{0})-u_{0}\|_{B^{s}_{p,r}}\geq ct^{1-\alpha}n^{-2}2^{{n}}-Ct^{1-\alpha}-C(2^{{n}}t)^{s}t^{-\alpha},$
Thus, picking $t^{1-\alpha}_{n}=n^{3}2^{-n}$ with large $n$, we have
$t^{-\alpha}_{n}\|\mathbf{S}_{t_{n}}(u_{0})-u_{0}\|_{B^{s}_{p,r}}$ $\geq cn-Cn^{3}2^{-n}-Cn^{3s}t_{n}^{\alpha(s-1)}\geq\tilde{c}n.$
This completes the proof of Theorem 1.1.
3.4 Proof of Theorem 1.2
Following the similar argument as in Section 3.2, we can establish the following Proposition.
Proposition 3.2.
Assume that $v_{0}=u_{0}$ satisfies (3.9). Under the assumptions of Theorem 1.1, we have
$\|\mathbf{S}_{t}\left(v_{0}\right)-v_{0}\|_{B^{s-1}_{p,r}}\lesssim t.$ (3.20)
Furthermore, there holds
  • for $s>\max\left\{1+\frac{1}{p},\frac{3}{2}\right\}$ and $(p,r)\in[1,\infty]\times[1,\infty),$ we have
    $\|\mathbf{\widetilde{w}}\|_{B^{s-2}_{p,r}}\lesssim t^{2};$
    here and in what follows we denote
[…]
  • $1+\frac{1}{p}<s\leq\frac{3}{2}$ and $(p,r)\in[1,\infty]\times[1,\infty)$ or $s=1+\frac{1}{p}$ and $(p,r)\in[1,\infty)\times\{1\}$, we have
    $\|\mathbf{\widetilde{w}}\|_{B^{0}_{p,r}}\lesssim t^{s}.$
Now we present the proof of Theorem 1.2.
Case 1: $s>\max\left\{1+\frac{1}{p},\frac{3}{2}\right\}$ and $(p,r)\in[1,\infty]\times[1,\infty)$.
By the triangle inequality and Propositions 3.2, we deduce that
$\|\mathbf{S}_{t}(v_{0})-v_{0}\|_{B^{s}_{p,r}}$ $\geq 2^{{ns}}\big{\|}\Delta_{n}\big{(}\mathbf{S}_{t}(v_{0})-v_{0}\big{)}\big{\|}_{L^{p}}=2^{{ns}}\big{\|}\Delta_{n}\big{(}t\mathbf{\widetilde{v}}_{0}+\mathbf{\widetilde{w}}\big{)}\big{\|}_{L^{p}}$
$\geq t2^{{ns}}\|\Delta_{n}\mathbf{\widetilde{v}}_{0}\|_{L^{p}}-2^{{2n}}\|\mathbf{\widetilde{w}}\|_{B^{s-2}_{p,\infty}}$
$\geq t2^{{n}s}\|v_{0}^{2}\partial_{x}\Delta_{n}v_{0}\|_{L^{p}}-t2^{{n}s}\|[\Delta_{n},v_{0}^{2}]\partial_{x}v_{0}\|_{L^{p}}$
$\quad-Ct\|\mathbf{Q}(v_{0})\|_{B^{s}_{p,\infty}}-C2^{2{n}}t^{2}$
$\geq ctn^{-2}2^{n}-Ct-C2^{2{n}}t^{2}.$
Case 2: $1+\frac{1}{p}<s\leq\frac{3}{2}$ and $(p,r)\in[1,\infty]\times[1,\infty)$ or $s=1+\frac{1}{p}$ and $(p,r)\in[1,\infty)\times\{1\}$.
By the triangle inequality and Propositions 3.2, we deduce that
$\|\mathbf{S}_{t}(v_{0})-v_{0}\|_{B^{s}_{p,r}}$ $\geq 2^{{ns}}\big{\|}\Delta_{n}\big{(}\mathbf{S}_{t}(v_{0})-v_{0}\big{)}\big{\|}_{L^{p}}=2^{{ns}}\big{\|}\Delta_{n}\big{(}t\mathbf{\widetilde{v}}_{0}+\mathbf{\widetilde{w}}\big{)}\big{\|}_{L^{p}}$
$\geq t2^{{ns}}\|\Delta_{n}\mathbf{\widetilde{v}}_{0}\|_{L^{p}}-2^{{ns}}\|\mathbf{\widetilde{w}}\|_{B^{0}_{p,r}}$
$\geq ctn^{-2}2^{n}-Ct-C(2^{{n}}t)^{s}.$
The remaining process can be done as above, we omit it. Thus we complete the proof of Theorem 1.2.
Acknowledgments
J. Li is supported by the National Natural Science Foundation of China (12161004), Training Program for Academic and Technical Leaders of Major Disciplines in Ganpo Juncai Support Program(20232BCJ23009) and Jiangxi Provincial Natural Science Foundation (20224BAB201008). Y. Yu is supported by National Natural Science Foundation of China (12101011). W. Zhu is supported by National Natural Science Foundation of China (12201118) and Guangdong Basic and Applied Basic Research Foundation (2021A1515111018).
Declarations
Data Availability No data was used for the research described in the article.
Conflict of interest The authors declare that they have no conflict of interest.